HELP!!! PLEASE!!
Solve: sqr is for square root.
7((2x-5)-(x+1))=(sqr7 +2)(sqr7 -2)

Answer by stanbon(75887)

Solve: sqr is for square root.
7((2x-5)-(x+1))=(sqr7 +2)(sqr7 -2)
7(x-6) = 7-4
7x-42 = 3
7x = 45
x = 45/7
===============
Cheers,
Stan H.